OpenStudy (anonymous):

a card is selected at random from a standard deck of 52 playing cards. Find the probability that the card is a 2, 3, 4, or 5 given that it is black. A.) 1/26 B.) 1/13 C.) 2/3 D.) 4/13 i got D but im not sure plz help

OpenStudy (anonymous):

youre right its D

OpenStudy (anonymous):

okay thanks what makes you sure?

OpenStudy (anonymous):

i dont thing the 'black' part has anything to do with it, but we can compute anyway. there are 26 black cards and of those 8 are 2,3,4, or 5 so it is \[\frac{8}{26}=\frac{4}{13}\]

OpenStudy (anonymous):

there are 8 total cards in question (4 cards x 2 per card(black)) out of 26 black cards total..8/26 = 4/13

OpenStudy (anonymous):

alright thanks for your help! :)